Which of the following compounds has the highest lattice energy? A) CaO B) CsI C) BaS D) NaF E) NaCl
Answer #1

lattice energy = strongest bonding, i.e. strongest electrongativity difference

A)

CaO = 1.3-3.5 = -2.2

CsI = 0.8 - 2.7 = -1.9

BaS = 0.9-2.5 = -1.6

NaF = 0.9 - 4 0 -3.1

NaCl = 0.9 - 3= -2.1

strongest lattice if NaF